Many doors in public places are fitted with simple and inexpensive mechanisms that are designed to close the doors while reduM = 70kg C = 1Ns/m k = 26N/m FO = 5N/S 9 = 95 (duration of the couple in seconds) b = 45N (magnitude of the step in N) z = 2.
Answer #1

mer R XE р - Apply Laplace 390 = (m84+k++8)x) TIF = 5 708² X+ 827 sau AS (3) 70²18+26 V> 5+4-3) (70148426) AS(S+2+4) 5 Х 8 +4dk 20 H KS (8) (5+43) fs+S426) KS (3424) = -($083430317 20:38+9+ K= -(7083+30387 30:38+)118) sull} 9 V S (S+4) dk (84445) FBCE >> (8423(708248 +26) +5 (+24) = 708°+ 30483+3038+11.8 +63 +2:46 29 7083 +(3024k +(302 4k]g?+ 30:38 3 +111.8 tdk 70 53 70 3